A simple circuit is constructed in which a 1 Ω resistor is connected across a 1 V battery and so draws a 1 A current. What is the current drawn from the battery if:

(a) an extra 1 Ω resistor is connected in series with the existing one?

(b) four extra 1 Ω resistors are connected in series with the existing one?


1
Expert's answer
2021-01-05T12:02:50-0500
  1. "I=\frac{U}{R}=\frac{1}{1+1}=0.5A"
  2. "I=\frac{U}{R}=\frac{1}{5\times1}=0.2A"